the largest root of an equation (x-1)(x-1)=4x-7
Answer by xcentaur(357)

(x-1)(x-1)=4x-7
(x-1)^2=4x-7
x^2+1-2x=4x-7
x^2+1-2x-4x+7=0
x^2-6x+8=0
(x-4)(x-2)=0
x=4,2
Larger root is 4
That is your answer


Hope this helps,
Prabhat